OpenStudy (anonymous):

What kind of triangle is formed if the side lengths are 12, 16, and 20? A. Acute B. Obtuse C. Right D. No triangle can be formed with those side lengths.

OpenStudy (zpupster):

I know this because dividing those given sides by 4 will give 3,4,5 and that is a right triangle by Pythagorean theorem a^2 + b^2 = c^2 or plugging them in 12^2 + 16^2 = 20^2 144 + 256 = 400 and this checks I will leave disproving whether it is acute, obtuse or none up to you.